OpenStudy (tux):

Given a group of 10 men and 5 women: a) how many ways can you choose a committee with 4 members b) How many combinations of 4 have 2 men on them c) have less men than women d) if Mary and John won't serve together how many 4 members combinations are possible?

OpenStudy (anonymous):

a) \[C _{4}^{15}=1365\]

OpenStudy (anonymous):

b) \[C_{2}^{10}\times C _{2}^{5}=450\]

OpenStudy (anonymous):

c) \[C _{1}^{10}\times C _{3}^{5}=100\]

OpenStudy (anonymous):

d)\[2\times C _{4}^{14}- C _{4}^{13}=1287\]